• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name flay

    Log entry time 07:58:28 on December 16, 2009

    Entry number 307413

    keyword=12/16/09 Owl Shift Summary

    00:00 Shift begins in production mode during run set 14698/26286/5792

    00:45 Dithering not working after EPICS reboot -- Diancheng working on it

    01:20 Diancheng decides to continue production running without dithering -- Will fix problem later in evening/day

    02:14 Diancheng fixed the dithering problem. See HALOG# 307372

    02:47 Beam has been down for ~45 mins. MCC checking Cryo Modules

    03:55 Beam still down

    06:09 Beam still down

    06:30 MCC has brought beam back

    07:00 Resuming production data taking

    Runs (PVDIS/LHRS/RHRS)

    14698/26286/5792 production - 38.4/488.6k good pairs 14699/26287/5793 production - 8.70/497.3k good pairs 14700/26288/5794 production - 7.90/505.2k good pairs 14701/26289/5795 production - 40.3/545.5k good pairs 14702/ production - 2.30/547.8k good pairs 14703/26290/5796 cosmics 14704/26291/5797 cosmics 14705/26292/5798 cosmics /26293/5799 spot++ 14706/26294/5800 production -